through a cross section of discharge tube current strength is 5 amphere. if the same no of negative electron and proton are crossing per section then calculate the number of negative electron crossing per section through crossing per section through cross section
1
Expert's answer
2018-09-06T16:41:08-0400
An electric current I is a flow of electric charge Q per t seconds:
I=Q/t
. In electric circuits the electric charge is mostly presented by electrons – negatively charged particles. The direction of the current is supposed to be the direction of positively charged particles. In the discharge tube negative electrons and positive protons move in the opposite directions. It means that the whole current 5 A consists of two currents: flow of electrons (for instance, they move left, and the half-current is directed right) and flow of protons (they move right, and the part-current is directed right). The number of electrons and protons is the same, so calculate the current created only by electrons:
I_e=I/2.
The number of electrons (charge of one electron is e=1.6·〖10〗^(-19) C) crossing the cross-section per one second:
N=Q_e/e=(I_e·t)/e=(I·t)/2e=(5·1)/(2·1.6·〖10〗^(-19) )=1.56·〖10〗^19